OpenStudy (anonymous):

Find the center of the ellipse with the equation 4x^2+3y^2-32x+18y-5=0

OpenStudy (mertsj):

\[4(x^2-8x)+3(y^2+6y)=5\]

OpenStudy (mertsj):

Complete the square twice. Once for x and once for y

OpenStudy (anonymous):

(3,-4) ?

OpenStudy (mertsj):

\[4(x^2-8x+16)+3(y^2+6y+9)=5+64+27\]

OpenStudy (mertsj):

(4,-3)